Determine the atmospheric pressure and calculate the solubility of oxygen in the lake in the Rocky Mountains. The Henry's law constant at 5°C is 1.6 x10-3 mol/lit atm. The total atmospheric pressure in Denver is 0.66 atm. The abundance of O2 remains constant at 20.95%.
Answer #1

Mole fraction of O2 = Volume fraction = %O2 in air / 100 = 0.2095

From Henry's law, we have the relation:

PO2 = mole fractionn of O2 * Atmospheric pressure = 0.2095 * 0.66 = 0.1383 atm

Solubility = KH * PO2 = 1.6 * 10-3 * 0.1383 = 2.213 * 10-4 mol/liter